OpenStudy (anonymous):

x^6 + y^6 = 1 What does: y'' = My book says y'' = (-5x^4) / (y^11) But I don't understand how. My work so far: y' = 6x^(5) + 6y^(5) * dy/dx = 0 y' = (-x^5)/(y^5) Now I use quotient rule. (f'g - fg')/(g^2) y'' = [(-5x^4)(y^5) + (x^5)(5y^4)(dy/dx)] / (y^10) Then i substituted (-x^5)/(y^5) for (dy/dx) Then I simplified, and factored. Ended up with (25x^10 - 5x^4) / (y^10)

OpenStudy (anonymous):

If you need me to show more work let me know.

OpenStudy (anonymous):

I've been watching the patrickJMT Calcus videos and they are pretty helpful.

OpenStudy (anonymous):

Ima double check.

OpenStudy (mathteacher1729):

Hold on I'm gonna check this out. (getting pencil & paper)

OpenStudy (anonymous):

yeah... your second derivative is incorrect...

OpenStudy (anonymous):

this is what I got: (-5x^10 - 5x^4*y^6)/y^11

OpenStudy (anonymous):

when you factor out (-5x^4) from that you'll end up with (x^6 + y^6) = 1

OpenStudy (eyust707):

just solve for y and take two derivatives

OpenStudy (eyust707):

there may be a point in there where you canepress a term in terms of y

OpenStudy (anonymous):

Wolfram doesn't do implicit differentiation to the best of my knowledge.

OpenStudy (eyust707):

ohhh haha im a dummy yea that way woud be way easier...

OpenStudy (anonymous):

dplanc how did you get that.

OpenStudy (anonymous):

Yeah I usually use wolfram. It's pretty good.

OpenStudy (anonymous):

i just did exactly what you did... double check your second derivative... btw... instead of using quotient rule, I used product rule: (-x^5)*(y^-5)

OpenStudy (anonymous):

kk

OpenStudy (mathteacher1729):

Ok, I've done this problem twice now (double-checking each method), once using the quotient rule the other using the product rule to differentiate -x^5/y^5 (or (-x^5*y^(-5) if you like). For both methods I obtain: (-5x^4y^6 -5x^10)/y^11

OpenStudy (mathteacher1729):

Something happened: y'' = [(-5x^4)(y^5) + (x^5)(5y^4)(dy/dx)] / (y^10) Then i substituted (-x^5)/(y^5) for (dy/dx) Then I simplified, and factored. <<<=== RIGHT HERE :-p Either that or it is way past time for me to be doing math and I should be going to sleep.

OpenStudy (anonymous):

ok =P ty I'ma stare really hard at this until I get it TT.

OpenStudy (mathteacher1729):

So let's take take the second term in the numerator: +5x^5*y^4*(-x^5/y^5) = -x^10/y

OpenStudy (anonymous):

Ok i get that one

OpenStudy (mathteacher1729):

Oh my goodness. I typed it all up and it had errors: This version is ok. You should have the fraction simplified like this, multiply by y/y to remove the irritating y from the numerator's fraction... \[\huge \frac{-5x^4y^5-\frac{5x^{10}}{y}}{y^{10}}\cdot\frac{y}{y}\] And this simplifies to \[\huge \frac{-5x^4y^6-5x^{10}}{y^{11}}\] And that's it. I mean you can factor out a -5x^4 if you like.

OpenStudy (anonymous):

Yeah I got that too at first but I thought I did something terribly wrong because my book definitely gives the wrong answer TT. Thank you very much for all the help =D.

OpenStudy (mathteacher1729):

No prob. :) Have a good evening!